Suppose a perfectly competitive firm with total cost function given as:
TC= 400+20Q-2Q2+2
3Q
3.
A. Find profit maximizing level of output and the maximum profit if the average revenue
equals $180.
B. Calculate the shutdown level of output and price.
1
Expert's answer
2021-08-10T12:14:42-0400

"TC=400 + 20Q - 2Q^2 + 23Q^3"

A. For a perfectly competitive firm average revenue is also equal to marginal revenue and thus price.

"P= AR = \$180"

Equilibrium condition in a perfectly competitive market;

"P = MC"

Marginal cost,

"MC = \frac{dTC}{dQ}"

"MC = 20 - 4Q + 69Q^2"

So,

"180 = 20 - 4Q + 69Q^2\\

69Q^2 - 4Q - 160 = 0"

Solving for Q

"Q = 1.5"

B. The shutdown rule states that in the short run a firm should continue to operate if price exceeds average variable costs.

"VC = 20Q - 2Q^2 + 23Q^3\\



AVC = \frac{VC}{Q}\\

AVC = 20 - 2Q + 23Q^2"

So, shut down level will be where;

"P = AVC\\

180 = 20 - 2Q + 23Q^2\\

23Q^2 - 2Q - 160 = 0\\"

Solving for Q

"Q = 2.7\\



P = 180"